# A Heuristic Philosophical Discourse on Various Applications of

### v1

@@ 6.1 The Issue of Functoriality
In [80] it has been shown that the aforesaid PGC of GR, which is tantamount to $\mathbf{A}$-invariance in ADG-gravity, is an example of the Functoriality of ADG-gravity. In other words, since the dynamics is categorically represented as equations involving the connection sheaf morphism $\mathcal{D}$ of the ADG-Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$, the PGC is represented by functors that preserve the relevant categories. [^26] In category-theoretic parlance, a functor between two categories, is a map or transformation that respects the objects and arrows of the two categories. In [80], in continuation and extension of [58, 59, 60], it was shown and argued that the relevant functors are, in fact, special types of $\mathbf{A}$-preserving functors (or $\mathbf{A}$-morphisms) called geometric morphisms, which preserve the ‘geometric’ structure of the vector sheaves involved in the dynamics.
@@
Aphorism 7. In ADG-Gravity, the dynamics is purely gauge and background spacetime manifold independent and functorial, while the PGC is functorially represented in terms of the principal group sheaf of automorphisms of the relevant vector sheaves as $\mathbf{A}$-invariance
citemall14,mall15,mall13,rap15. The latter simply means that the ADG-field dynamics, which in its Einstein-Hilbert action expression involves the curvature of the connection which is an $\otimes_{\mathbf{A}}$-tensor, remains invariant (or ‘unperturbed’) by our generalised coordinate ‘measurements’ that are organised in the structure algebra sheaf $\mathbf{A}$.(E1,M1,F1,E2,M2)
Thus, all our generalised measurements are represented in ADG as sections of the structure algebra sheaf $\mathbf{A}$ [^27] The reader should note here that the structure sheaf $\mathbf{A}$ is supposed to be a sheaf of abelian (:commutative) algebras. This reflects our primitive assumption that we always measure commutative numbers, while in the Quantum Theory it is supposed to be the ADG-version of Bohr’s Correspondence principle: although quantum observables may be noncommutative q-numbers, our measurements thereof are commutative c-numbers. In turn, the ADG-gravitational dynamics, since it is $\mathbf{A}$-functorial, [^28] Or as Mallios originally coined it: $\mathbf{A}$-invariant. is not ‘disturbed’ at all by our generalised field measurements in $\mathbf{A}$. Furthermore, the $\otimes_{\mathbf{A}}$-functorial ADG-gravitational field dynamics does not break down in any differential geometrical sense in the presence of any type of singularity that may be encoded in the structure sheaf $\mathbf{A}$.
@@ 7 Gauge Field Theory of the Third Kind and its Third Quantisation
$\bullet$ Gauge Field Theory of the Third Kind. We discussed earlier how from an ADG-theoretic perspective gravity is regarded as a gauge theory. We noted that the ADG-formalism may be coined half-order formalism, to distinguish it from the original second-order formalism of Einstein, whereby the dynamical variable is the smooth Riemannan spacetime metric $g_{\mu\nu}$, or from the more recent first-order formalism of Ashtekar, whereby the gravitational dynamical variables are the spin-connection $\mathcal{A}_{\mu}$ and the vierbein frame $e_{\mu}$. In ADG-gravity, the sole dynamical variable is the Einstein (connection) field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$.
@@
The second denomination gauge theory of the third kind comes from the observation that the first $U(1)$ gauge (or scale) theory for the electromagnetic and the gravitational field due to Weyl [103] was a global gauge theory, [^29] In the sense that Weyl showed that non-spacetime localised (global) gauge/scale invariance implies the conservation of electric charge in much the same way that general coordinate invariance leads to the conservation of energy and momentum in gravitational dynamics. while the gauge theories underlying the three fundamental forces (other than gravity) of the Standard Model [^30] That is, the electromagnetic (with local gauge group $U(1)$), the weak nuclear (with local gauge group $SU(2)$) and the strong nuclear (with local gauge group $SU(3)$) forces. are flat Minkowski space localised gauge theories [24].
By contrast, aside from its half-order formalism, our ADG-based gauge-theoretic formulation of gravity (:ADG-gravity), although local by its sheaf-theoretic character, is not background spacetime localised, since there is no background spacetime manifold to localise and solder it on to begin with.(E1,M1, F1,E2,M2)
$\bullet$ Third Quantisation. In [78, 79], and recently in [80], a third canonical type of ADG-field quantisation scenario was proposed according to which certain local, sheaf cohomological characteristic forms for both the vector sheaf part $\mathbf{\mathcal{E}}$ and the connection part $\mathcal{D}$ of the ADG-theoretic vacuum Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$ were seen to obey canonical type of Heisenberg non-commutation relations, albeit, explicitly not parametrised by a background spacetime manifold. [^31] In the sense that they are not equal-time commutation relations which, in the usual canonical QG scenario, would have been required to obey some global hyperbolicity type of foliation of the background spacetime manifold into time-parametrised 3-dimensional spacelike hypersurfaces. This is to be expected as our ADG-gravity does not depend at all on an external (background) spacetime manifold, as well as to be desired, as our third quantisation scenario would be expected to ‘algebraically close’ within the autonomous and ‘solipsistic’ ADG-theoretic Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$.
@@
8 Brief Philosophical Musings on the Future of ADG-Gravity

### v2

@@ 6.1 The Issue of Functoriality
In [85] it has been shown that the aforesaid PGC of GR, which is tantamount to $\mathbf{A}$-invariance in ADG-gravity, is an example of the Functoriality of ADG-gravity. In other words, since the dynamics is categorically represented as equations involving the connection sheaf morphism $\mathcal{D}$ of the ADG-Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$, the PGC is represented by functors that preserve the relevant categories. [^29] In category-theoretic parlance, a functor between two categories, is a map or transformation that respects the objects and arrows of the two categories. In [85], in continuation and extension of [63, 64, 65], it was shown and argued that the relevant functors are, in fact, special types of $\mathbf{A}$-preserving functors (or $\mathbf{A}$-morphisms) called geometric morphisms, which preserve the ‘geometric’ structure of the vector sheaves involved in the ADG-gravitational field autodynamics.
@@
Aphorism 7. In ADG-Gravity, the dynamics is purely gauge and background spacetime manifold independent and functorial, while the PGC is functorially represented in terms of the principal group sheaf of automorphisms of the relevant vector sheaves as $\mathbf{A}$-invariance [63, 64, 65, 85]. The latter simply means that the ADG-field dynamics, which in its Einstein-Hilbert action expression involves the curvature of the connection which is an $\otimes_{\mathbf{A}}$-tensor, remains invariant (or ‘unperturbed’) by our generalised coordinate ‘measurements’ that are organised in the structure algebra sheaf $\mathbf{A}$.(E1,M1,F1,E2,M2)
Thus, all our generalised measurements are represented in ADG as sections of the structure algebra sheaf $\mathbf{A}$ [^30] The reader should note here that the structure sheaf $\mathbf{A}$ is supposed to be a sheaf of abelian (:commutative) algebras. This reflects our primitive assumption that we always measure commutative numbers (:Bohr’s original $c$-numbers), while in the Quantum Theory it is supposed to be the ADG-version of Bohr’s Correspondence principle: although quantum observables may be noncommutative $q$-numbers, our measurements thereof are commutative $c$-numbers. In turn, the ADG-gravitational autodynamics, since it is $\mathbf{A}$-functorial, [^31] Or as Mallios originally coined it: $\mathbf{A}$-invariant. is not ‘disturbed’ at all by our generalised field measurements in $\mathbf{A}$. Furthermore, the $\otimes_{\mathbf{A}}$-functorial ADG-gravitational field dynamics does not break down in any differential geometrical sense in the presence of any type of singularity that may be encoded in the structure sheaf $\mathbf{A}$ of our generalised coordinates (:arithmetics).
@@ 7 Gauge Field Theory of the Third Kind and its Third Quantisation
$\bullet$ Gauge Field Theory of the Third Kind. We discussed earlier how from an ADG-theoretic perspective gravity is regarded as a gauge theory. We noted that the ADG-formalism may be coined half-order formalism, to distinguish it from the original second-order formalism of Einstein, whereby the dynamical variable is the smooth Riemannan spacetime metric $g_{\mu\nu}$, or from the more recent first-order formalism of Ashtekar, whereby the gravitational dynamical variables are the spin-connection $\mathcal{A}_{\mu}$ and the vierbein comoving frame $e_{\mu}$. In ADG-gravity, the sole dynamical variable is the Einstein (connection) field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$.
@@
The second denomination, gauge theory of the third kind, comes from the observation that the first $U(1)$ gauge (or scale) theory for the electromagnetic and the gravitational field due to Weyl [107] was a global gauge theory, [^32] In the sense that Weyl showed that non-spacetime localised (global) gauge/scale invariance implies the conservation of electric charge in much the same way that general coordinate invariance leads to the conservation of energy and momentum in gravitational dynamics. Weyl’s original gauge theory is commonly referred to as gauge theory of the 1st kind. while the local gauge field theories underlying the three fundamental forces (other than gravity) of the Standard Model [^33] That is, the electromagnetic (with local gauge group $U(1)$), the weak nuclear (with local gauge group $SU(2)$) and the strong nuclear (with local gauge group $SU(3)$) forces. are flat Minkowski space localised gauge theories [23]. [^34] And they are commonly referred to as gauge theories of the 2nd kind.
By contrast, aside from its half-order formalism, our ADG-based gauge-theoretic formulation of gravity (:ADG-gravity), although local by its very sheaf-theoretic character, is not background spacetime localised, since there is no background spacetime manifold to localise and solder it on to begin with.(E1,M1, F1,E2,M2)
$\bullet$ Third Quantisation. In [83, 84], and recently in [85], a third canonical type of ADG-field quantisation scenario was proposed according to which certain local, sheaf cohomological characteristic forms that characterise both the vector sheaf part $\mathbf{\mathcal{E}}$ and the connection part $\mathcal{D}$ of the ADG-theoretic vacuum Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$ were seen to obey canonical (:generalised position-momentum) Heisenberg type commutation relations, albeit, explicitly not parametrised by a background spacetime manifold. [^35] In the sense that they are not equal-time commutation relations which, in the usual canonical QG scenario, would have been required to obey some global hyperbolicity type of foliation of the background spacetime manifold into time-parametrised 3-dimensional spacelike hypersurfaces. This is to be expected as our ADG-gravity does not depend at all on an external (background) spacetime manifold, as well as to be desired, as our third quantisation scenario would be expected to ‘algebraically close’ within the autonomous and ‘solipsistic’ ADG-theoretic Einstein field $\mathcal{F}_{Einst}=(\mathbf{\mathcal{E}},\mathcal{D})$.
@@
8 Very Brief Philosophical Musings on the Immediate Future of ADG-Gravity